# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v3

@@ A.3.2 Derivation for Variational Lower Bound
Since $q_{\phi}$ only involves linear transformations and a softmax function, we argue that minimizing the KL loss between $q_{\phi}^{t}$ and $q_{\phi}^{t+1}$ is equivalent to minimizing the difference between $\phi^{t}$ and $\phi^{t+1}$.
Firstly, the logits $s_{i}$ are computed as: $s_{i}=\mathbf{W}_{k}^{(l)}\mathbf{e}^{k}$. A change in $\mathbf{W}_{k}^{(l)}$ , denoted as $\Delta\mathbf{W}_{k}^{(l)}$ , modifies $s_{i}$ as $\Delta s_{i}=(\Delta\mathbf{W}_{k}^{(l)})\mathbf{e}^{k}$. The change in $s_{i}$ is linear with respect to $\Delta\mathbf{W}_{k}^{(l)}$ and $\mathbf{e}^{k}$ Therefore, a small change in $\mathbf{W}_{k}^{(l)}$ leads to proportionally small changes in $s_{i}$.
Secondly, the Softmax function introduces nonlinear coupling between logits $s_{i}$ , as the output probabilities $\pi_{i}$ depend not only on $s_{i}$ but also on all other logits $s_{j}$ For a small change in $s_{i}$ , the change in $\pi_{i}$ can be approximated using the gradient of the Softmax $\frac{\partial\pi_{i}}{\partial s_{i}}=\pi_{i}(1-\pi_{i})$, $\frac{\partial\pi_{i}}{\partial s_{j}}=-\pi_{i}\pi_{j}\quad\text{for }i\neq j$. Thus, a change in $\mathbf{W}k^{(l)}$ affects $\pi_{i}$ as $\Delta\pi_{i}\approx\pi_{i}(1-\pi_{i})\Delta s_{i}-\sum_{j\neq i}\pi_{i}\pi_{j}\Delta s_{j}$. This implies that the change of the output $\Delta\pi_{i}$ shrinks proportionally with the change of parameters $\Delta\mathbf{W}_{k}$.
Lastly, for small parameter changes, we can approximate the KL divergence between the two categorical distributions using a second-order Taylor expansion around $q_{\phi}^{t}$. The KL divergence is defined as:
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)=\sum_{i}\pi_{i}^{t}\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right).$ (12)
Expanding $\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right)$ around $\pi_{i}^{t}$ using the Taylor series for $\log(1+x)$ where $x=-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}$ and keeping terms up to second order, we obtain:
$\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right)$ $=\log\left(1+\frac{\pi_{i}^{t}-\pi_{i}^{t+1}}{\pi_{i}^{t}}\right)=\log\left(1-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)\approx-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}.$ (13)
Substituting this into the KL divergence expression and ignoring higher-order terms (since $\Delta\pi_{i}$ is small), we get:
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)$ $\approx\sum_{i}\pi_{i}^{t}\left(-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}\right)=-\sum_{i}\Delta\pi_{i}-\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}.$ (14)
Since $\sum_{i}\Delta\pi_{i}=0$ (as probabilities sum to one), the first term vanishes, leaving:
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)\approx\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}.$ (15)
Substituting the expression for $\Delta\pi_{i}$ from above and noting that $\Delta s_{i}$ is linear in $\Delta\phi$, we observe that the KL divergence is a quadratic function of $\Delta\phi$. Therefore, for small $\Delta\phi$, the KL divergence can be approximated as:
$\mathcal{L}_{\text{KL}}\approx\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}\propto\|\Delta\phi\|^{2}=\|\phi^{t}-\phi^{t+1}\|^{2}.$ (16)
Thus, minimizing the KL divergence is approximately equivalent to minimizing $\|\phi^{t}-\phi^{t+1}\|^{2}$. In our setting, we use weight decay to regularize the model, which indirectly helps control the KL loss. Weight decay adds an L2 penalty to the loss function that encourages smaller parameter values, effectively shrinking the magnitude of the weights during training. The update rule for parameters with weight decay is given by: $\phi_{t+1}=\phi_{t}-\alpha\cdot\nabla_{\phi}\mathcal{L}_{\text{original}}-\alpha\cdot\lambda\phi_{t}$, where $\alpha$ is the learning rate and $\lambda$ is the weight decay coefficient. The total KL loss can be approximated as:
$\mathcal{L}_{\text{kl}}\approx\|\phi^{t}-\phi^{t+1}\|^{2}=\alpha^{2}\|\nabla_{\phi}\mathcal{L}_{\text{original}}+\lambda\phi_{t}\|^{2}.$ (17)
When the training process converges to a minimum, the task-related gradients ($\nabla_{\phi}\mathcal{L}_{\text{original}}$) become small and nearly zero, and the KL loss becomes dominated by the weight decay term, and the approximation simplifies to: $\mathcal{L}_{\text{kl}}\approx\alpha^{2}\lambda^{2}\|\phi^{t}\|^{2}$. Thus, the L2 norm contributes to minimizing the KL loss. ∎
Next, we provide proof of the theorem for the M-Step.
Theorem A.2.
Assuming $q^{t+1}_{\phi_{t}}(Z)=p^{t+1}(Z)$ and an L2 norm is applied on $\phi$, the variational lower bound in Eq. (3) can be approximated as follows:
$\mathbb{E}_{p(X)}\left[\mathbb{E}_{q^{t+1}_{\phi_{t}}(Z|X)}\left[\log p^{t+1}_{\Theta_{t+1}}(\hat{Y}|X,Z)\right]\right]-C,$ (18)
which is equivalent to minimizing the expected reconstruction loss $\mathbb{E}_{\mathbf{x}\sim p(X)}[\mathcal{L}_{recon}(\mathbf{x},\hat{\mathbf{x}})]$
@@
In the M-step, $q_{\phi}(Z|X)$ serves as the estimator for the environment distribution $p^{t+1}(Z)$ defined by $q^{t+1}_{\phi_{t}}(Z)$ in the previous E-step. Then, we derive the expected log-likelihood for $\log p_{\Theta}(\hat{Y}|do(X))$ from Eq. (3) :
$\quad E_{p(X)}\log{p_{\Theta}(\hat{Y}|do(X))}$
$=\sum_{\mathbf{x}}\log{p_{\Theta}(\hat{Y}|do(X)=\mathbf{x})}p(X=\mathbf{x})$
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}p(Z=z)$
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}p(Z=z)\cdot\frac{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$\geq\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
(Jensen’s Inequality)
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}$
$\hskip 58.32823pt+\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
$=\underbrace{\mathbb{E}_{p(X)}\mathbb{E}_{q^{t+1}_{\phi_{t}}(Z|X)}\log{p_{\Theta}(\hat{Y}|X,Z)}}_{-\mathcal{L}_{recon}}$
$+\underbrace{\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}}_{-\mathcal{L}_{reg}}.$
The first term, similar to the justification in the E-step, maximizes the predictive power of the model in a batch-wise manner, while the second term serves as a regularization of the environment estimator. Since we assume $q^{t+1}_{\phi_{t}}(Z)=p(Z)$, we prove that the second term can be further reduced:
$-\mathcal{L}_{reg}$ (19)
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\log{p(Z=z)}$
$\hskip 28.45274pt-\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$=\sum_{z}\log{p(Z=z)}\sum_{\mathbf{x}}p(X=\mathbf{x})q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})+H_{\phi_{t}}(Z|X)$
$=\sum_{z}\log{q^{t+1}_{\phi_{t}}(Z=z)}q^{t+1}_{\phi_{t}}(Z=z)+H_{\phi_{t}}(Z|X)$
$=-H_{\phi_{t}}(Z)+H_{\phi_{t}}(Z|X)$
$=-I_{\phi_{t}}(Z;X),$
where $H_{\phi_{t}}(Z)$ and $H_{\phi_{t}}(Z|X)$ are the entropy and the conditional entropy induced by the estimator $q^{t+1}_{\phi_{t}}$ respectively, and $I_{\phi_{t}}$ denotes the mutual information. Since $I_{\phi^{t+1}}(Z;X)$ is approximated to a constant $C$ as $p_{s}(X)\approx p(X)$, we are able to acquire the final lower bound as:
$E_{p(X)}\log{p_{\Theta}(\hat{Y}|do(X))}\geq\mathbb{E}_{p(X)}\mathbb{E}_{q^{t+1}_{\phi}(Z|X)}\log{p_{\Theta}(\hat{Y}|X,Z)}-C$ (20)
Therefore we only focus on maximizing the first term, which becomes minimizing the reconstruction loss $\mathcal{L}_{recon}$. In practice, we make $q^{t+1}_{\phi}$ active during this step while posing an L2 norm on its parameters to better optimize the environment estimator as well as switching to a different environment distribution space. During training, as proven in Proof A.3.3, the updated $q^{t+1}_{\phi_{t+1}}$ eventually approximates the $q^{t+1}_{\phi_{t}}$.
A.4 Pre-train and Downstream Datasets Details
@@
We apply z-score normalization for all datasets to align disease data with zero-mean.
A.4.1 Pre-train Datasets
In this study, we utilize a comprehensive collection of 17 distinct diseases from the United States, sourced from Project Tycho. These diseases encompass both respiratory and non-respiratory categories and serve as the foundation for pre-training three transformer-based models: CAPE, PEM, and PatchTST. The selection criteria for these datasets were meticulously chosen based on the following factors:
@@ A.4.1 Pre-train Datasets
Table 6: Pre-training Datasets collected from Project Tycho.
Disease Number of States Total Length Non-Respiratory
Gonorrhea 39 37,824 Yes
Meningococcal Meningitis 37 44,890 No
Varicella 30 33,298 No
Typhoid Fever 44 89,868 Yes
Acute Poliomyelitis 47 74,070 Yes
Hepatitis B 31 34,322 Yes
Pneumonia 41 68,408 No
Hepatitis A 38 37,303 Yes
Influenza 42 61,622 No
Scarlet Fever 48 129,460 No
Smallpox 44 71,790 No
Tuberculosis 39 95,564 No
Measles 50 151,867 No
Diphtheria 46 112,037 No
Mumps 41 50,215 No
Pertussis 46 109,761 No
Rubella 7 6,274 No
A.4.2 Downstream Datasets
@@
Disease Number of Regions Sampling Rate Respiratory Total Length
ILI USA 1 Weekly Yes 966
ILI Japan 1 Weekly Yes 348
Measles 1 Bi-weekly Yes 1,108
Dengue 23 Mixed No 10,739
RSV 13 Weekly Yes 4,316
MPox 1 Daily No 876
COVID 16 Daily Yes 12,800
@@
  3. Measles  (Lau et al., 2020a): The measles dataset contains biweekly measles infections in England from 1906 to 1948.
@@ A.5 Implementation Details
  3. Our framework is extensible to spatiotemporal modeling. Specifically, the temporal input can be replaced with graph structured data, and the predictor can incorporate graph-based encoders. Exploring this direction is exciting future work, but we believe temporal pre-training is a crucial first step toward that goal.
Zero-Shot. Once pre-trained, our CAPE framework can be directly utilized for zero-shot forecasting where the model remains frozen and no parameter is updated. Similar to the MOMENT model (Goswami et al., 2024), we retain the pre-trained reconstruction head and mask the last patch of the input to perform forecasting: $\hat{\mathbf{y}}=\hat{\mathbf{x}}_{[T-c:T]}$.
@@
Model Details. We design our model by stacking 4 layers of the CAPE encoder, each with a hidden size of 512 and 4 attention heads. For environment representations, we incorporate 16 distinct environments, each encoded with a size of 512. To ensure a fair comparison, PatchTST is configured with the same number of layers and hidden size as our CAPE-based model. For all other baseline models, we adopt the architectures as reported in previous studies (Wang et al., 2024a; Kamarthi & Prakash, 2023; Panagopoulos et al., 2021).
@@
  1. We adopt an input length of 36 (Wu et al., 2023; Wang et al., 2024b) and a patch size of 4 for applicable models. For the environment estimator defined in Eq. (4), a shared weight $w_{k}$ is used for all environment representations. All results are evaluated using Mean Squared Error (MSE).
@@
  3. For the training process, we pre-train CAPE, PEM, and PatchTST on a single Nvidia A100 GPU. During pre-training, we utilize only 70% of the available training data, specifically the first 70% of the dataset for each disease category. We set the learning rate to $1\times 10^{-5}$ and apply a dropout rate of 0.1 to prevent overfitting. In the CAPE pre-training strategy, we assign a weight of 0.5 to $\alpha$ to balance the contributions of contrast loss to the whole loss function. A detailed illustration of our pre-training strategy is shown in Algorithm 1.
  4. After pre-training, we fine-tune the entire model using a single Nvidia K80 GPU, maintaining the same hyperparameter settings for consistency. The best-performing model is selected based on its performance on the validation set. Similarly, for all baseline models, we train each until convergence and select the optimal model based on validation set performance for the subsequent test.
Fine-Tuning. For downstream tasks, we fine-tune the entire model using MSE loss, still employing EM for optimization. The difference is that we replace the task-specific head $h_{\psi}$ for pre-training, which is a linear transformation on each patch, with a head for forecasting that takes the concatenated latent representations $\mathbf{x}^{(L)}$ and maps them to the future prediction: $\hat{\mathbf{y}}=h_{\psi}(\mathbf{x}^{(L)})=\mathbf{W}\mathbf{x}^{(L)}$.
Algorithm 1 Hierarchical Contrastive Loss Optimization with EM
1:Input: $\mathcal{D}_{tr}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, $\alpha$, $\eta_{\theta,\psi,\phi}$, $\eta_{\mathbf{E}}$, $T_{max}$
2:Initialize: $\theta$, $\psi$, $\phi$, $\mathbf{E}\sim\text{Orthogonal}(\mathbf{I})$
3:repeat
4:    E-Step: Optimize Environment Representations $\mathbf{E}$
5:     Freeze $(\theta,\psi)$ and set $\alpha\leftarrow 0$
6:     Sample mini-batch $\mathcal{B}\subseteq\mathcal{D}_{tr}$
7:     Compute reconstruction loss: $\mathcal{L}_{recon}$
8:     Compute gradient: $\nabla_{\mathbf{E}}\mathcal{L}_{recon}$
9:     Update $\mathbf{E}$: $\mathbf{E}\leftarrow\mathbf{E}-\eta_{\mathbf{E}}\nabla_{\mathbf{E}}\mathcal{L}_{recon}$
10:    M-Step: Optimize Model Parameters $(\theta,\psi,\phi)$
11:     Freeze $\mathbf{E}$, set $\alpha$ to predefined values
12:     Sample mini-batch $\mathcal{B}\subseteq\mathcal{D}_{tr}$
13:     Compute contrastive loss: $\mathcal{L}_{CL}$
14:     Compute total loss: $\mathcal{L}_{final}=\mathcal{L}_{recon}+\alpha\mathcal{L}_{CL}$
15:     Update parameters:
16:       $(\theta,\psi,\phi)\leftarrow(\theta,\psi,\phi)-\eta_{\theta,\psi,\phi}\nabla_{\theta,\psi,\phi}\mathcal{L}_{final}$
17:until $t=T_{max}$
18:Output: $(\theta,\psi,\phi),\mathbf{E}$
A.6 Evaluation Metrics
In this study, we primarily use MSE for its sensitivity to hard cases (large errors) and common use in epidemic forecasting  (Datilo et al., 2019). In addition, we also report MAPE and peak timing error (averaged over horizons 1–16) on the ILI USA dataset as an example:
Table 8: Model Performance Metrics
Model MAPE Peak Error
CAPE 0.188 1.226
Informer 0.246 1.553
LSTM 0.264 1.398
DLinear 0.225 1.442
A.7 Environment Influence
We also add Gaussian noise to environment weights (Equation 4), causing performance drops, which highlights the importance of accurate environment estimation:
@@ A.7 Environment Influence
A.8 Full Results on Pre-train datasets

### v4

@@ A.4.2 Proof of Equation  5
For any compatible $A,B$ and any unit vector $x$,
$\|ABx\|_{2}\;\geq\;\sigma_{\min}(A)\,\|Bx\|_{2}\;\geq\;\sigma_{\min}(A)\,\sigma_{\min}(B)\,\|x\|_{2},$
hence
$\sigma_{\min}(AB)\;\geq\;\sigma_{\min}(A)\,\sigma_{\min}(B).$
Apply (S1) with $A=F$ and $B=V^{-1}$ (with $V$ invertible):
$\sigma_{\min}(FV^{-1})\;\geq\;\sigma_{\min}(F)\,\sigma_{\min}(V^{-1}).$
Using the inversion identity for singular values,
$\sigma_{\min}(V^{-1})=\frac{1}{\sigma_{\max}(V)},$
we obtain
$\boxed{\;\sigma_{\min}(FV^{-1})\;\geq\;\dfrac{\sigma_{\min}(F)}{\sigma_{\max}(V)}.\;}$
By (P1) applied to $M=FV^{-1}$, we get:
$\sigma_{\min}(FV^{-1})\;\leq\;\rho(FV^{-1}).$
Combining (L1) and (L2) yields the rigorous lower bound
$\boxed{\;\dfrac{\sigma_{\min}(F)}{\sigma_{\max}(V)}\;\leq\;\rho(FV^{-1}).\;}$
Derivation for Upper bound.
@@
For any square $M$, $\rho(M)\leq\|M\|_{2}$ by (P2). Thus
$\rho(FV^{-1})\;\leq\;\|FV^{-1}\|_{2}\;\leq\;\|F\|_{2}\,\|V^{-1}\|_{2}=\sigma_{\max}(F)\,\frac{1}{\sigma_{\min}(V)}.$
Hence the rigorous upper bound is
$\boxed{\;\rho(FV^{-1})\;\leq\;\dfrac{\sigma_{\max}(F)}{\sigma_{\min}(V)}.\;}$
Final result. Putting (LB) and (UB) together, we obtain
$\boxed{\;\frac{\sigma_{\min}(F)}{\sigma_{\max}(V)}\;\leq\;\rho(FV^{-1})\;\leq\;\frac{\sigma_{\max}(F)}{\sigma_{\min}(V)}.\;}$
A.5 Pre-train and Downstream Datasets Details
@@
We collect the rough range of $R_{0}$ for each pathogen from various prior research and apply z-score normalization for all datasets.
A.5.1 Pre-train Datasets
In this study, we utilize a comprehensive collection of 17 distinct diseases from the United States, sourced from Project Tycho. These diseases encompass both respiratory and non-respiratory categories and serve as the foundation for pre-training two transformer-based models: CAPE, and PatchTST. The selection criteria for these datasets were meticulously chosen based on the following factors:
@@ A.5.1 Pre-train Datasets
Table 5: Pre-training datasets from Project Tycho with basic reproduction number ($R_{0}$) ranges.
Disease Number of States Total Length Non-Respiratory $R_{0}$ Range
Gonorrhea 39 37,824 Yes 1.00–1.01
Meningococcal Meningitis 37 44,890 No 0.6–1.6
Varicella 30 33,298 No 10–12
Typhoid Fever 44 89,868 Yes 2.8–7.0
Acute Poliomyelitis 47 74,070 Yes 5–7
Hepatitis B 31 34,322 Yes 1.0–3.3
Pneumonia 41 68,408 No 1.4–1.4
Hepatitis A 38 37,303 Yes 1.1–3.5
Influenza 42 61,622 No 1.2–1.6
Scarlet Fever 48 129,460 No 0.6–2.0
Smallpox 44 71,790 No 3.5–6
Tuberculosis 39 95,564 No 0.24–4.3
Measles 50 151,867 No 12–18
Diphtheria 46 112,037 No 1.7–4.3
Mumps 41 50,215 No 4–7
Pertussis 46 109,761 No 12–17
Rubella 7 6,274 No 3.4–7.0
A.5.2 Downstream Datasets
@@
Disease Number of Regions Sampling Rate Respiratory Total Length $\mathbf{R_{0}}$
ILI USA 1 Weekly Yes 966 1.2-1.4
ILI Japan 1 Weekly Yes 348 1.0-2.0
Measles 1 Bi-weekly Yes 1,108 12-18
Dengue 23 Mixed No 10,739 3.12-5.39
RSV 13 Weekly Yes 4,316 1-5
MPox 1 Daily No 876 1.1-2.7
COVID 16 Daily Yes 12,800 2.9-9.5
@@
  3. Measles  (Lau et al., 2020): The measles dataset contains biweekly measles infections in England from 1906 to 1948.
@@ A.6 Implementation Details
  3. Our framework is extensible to spatiotemporal modeling. Specifically, the temporal input can be replaced with graph-structured data, and the predictor can incorporate graph-based encoders. Exploring this direction is exciting future work, but we believe temporal pre-training is a crucial first step toward that goal.
Zero-Shot. Once pre-trained, our CAPE framework can be directly utilized for zero-shot forecasting, where the model remains frozen and no parameter is updated. Similar to the MOMENT model (Goswami et al., 2024), we retain the pre-trained reconstruction head and mask the last patch of the input to perform forecasting: $\hat{\mathbf{y}}=\hat{\mathbf{x}}_{[T-c:T]}$.
@@
Model Details. We design our model by stacking 4 layers of the CAPE encoder, each with a hidden size of 512 and 4 attention heads. For compartment representations, we incorporate 16 distinct compartments, each encoded with a size of 512. We constrain two prototypes with monotonic increase or decrease loss, respectively, 6 with the non-monotonic loss, and leave the rest to be unconstrained. To ensure a fair comparison, PatchTST is configured with the same number of layers and hidden size as our CAPE-based model. For all other baseline models, we adopt the architectures as reported in previous studies (Wang et al., 2024a; Kamarthi & Prakash, 2023; Panagopoulos et al., 2021).
@@
  1. We adopt an input length of 36 (Wu et al., 2023; Wang et al., 2024b) and a patch size of 4 for applicable models. For the compartment estimator defined in Eq. equation 1, a shared weight $w_{k}$ is used for all compartment representations. All results are evaluated using Mean Squared Error (MSE).
@@
  3. For the training process, we pre-train CAPE and PatchTST on a single Nvidia L40 GPU. During pre-training, we utilize only 70% of the available training data, specifically the first 70% of the dataset for each disease category. We set the learning rate to $1\times 10^{-5}$. In the CAPE pre-training strategy, we assign a weight of 1e-5 to $\lambda$ to balance the contribution of alignment loss to the whole loss function. We use a general $R_{0}$ range of 0-20 during pre-training.
  4. After pre-training, we fine-tune the entire model for five epochs with a changing learning rate, weight decay, and a larger $\lambda$. We also apply the disease-specific $R_{0}$ (shown in Table 6) for supervision during this phase. The best-performing model is selected based on its performance on the validation set. Similarly, for all baseline models, we train each until convergence and select the optimal model based on validation set performance for the subsequent test.
A.7 compartment Influence
We also add Gaussian noise to mixture weights (Equation 1), causing performance drops, which highlights the importance of accurate compartment estimation:
@@
A.8 Comparisons with PEM and ARIMA
In this section, we provide further comparisons with the Statistical model ARIMA (Panagopoulos et al., 2021), which is configured in an online forecasting setting (Pham et al., 2022; Abdulmajeed et al., 2020) where parameters are updated with each new sample. As shown in Table 8, ARIMA typically outperforms CAPE in the short-term forecasting, while CAPE outperforms ARIMA in long-term forecasting and the averaged performance.
Table 8: Mean Squared Error (MSE) comparison between ARIMA, CAPE, and PEM models. The lowest MSE for each horizon is marked in bold.
ILI USA ILI Japan Measles Dengue Covid
Horizon ARIMA CAPE PEM ARIMA CAPE PEM ARIMA CAPE PEM ARIMA CAPE PEM ARIMA CAPE PEM
1 0.138 0.174 0.303 0.358 0.328 0.734 0.070 0.111 0.330 0.244 0.367 0.912 33.779 25.841 36.163
2 0.203 0.192 0.328 0.772 0.709 0.919 0.120 0.157 0.350 0.373 0.317 0.844 33.199 25.413 29.278
4 0.354 0.299 0.507 1.720 1.191 1.310 0.223 0.188 0.464 0.696 0.508 1.236 32.476 24.631 33.545
8 0.702 0.469 0.519 2.991 1.792 1.836 0.481 0.406 0.726 1.736 1.169 1.806 36.567 33.003 39.577
16 1.119 0.650 0.682 2.590 1.878 1.936 1.047 0.883 1.213 4.131 2.512 2.938 42.908 49.838 49.299
Avg 0.503 0.357 0.468 1.686 1.179 1.347 0.388 0.349 0.616 1.436 0.975 1.547 35.785 31.745 37.573
A.9 Full Results on Pre-train datasets